$40,000 is invested at 5% interest. What is the amount of money in the account after 7 years if the interest is compounded annually?
Found 2 solutions by checkley77, penguin_129:
Answer by checkley77(12569)

40,000(1+.05)^7
40,000*1.05^7
40,000*1.4071=$56,284 ans.

Answer by penguin_129(5)

We use the formula for compound interest with P = 40,000 ; r = 5% = 0.05 ; m = 1 and t = 7.
The amount of the money in the account after 7 years:
A = P * ( 1 + r/m )^(t*m)

= 40,000 * ( 1 + .05/1 )^(7*1)

= $56,284